Find the value of the following expression for the given values.

n2−m2​; n=−3​, and m=-7


1
Expert's answer
2022-01-17T16:28:35-0500
"n^2-m^2=(-3)^2-(-7)^2=9-49=-40"

"n^2-m^2=(n-m)(n+m)"

"=(-3-(-7))(-3+(-7))=4(-10)=-40"